Find the arc length of the great circle on a sphere of radius 3000 miles subtended by a central angle of 7°. Express your answer in:
a. Degrees b. Nautical miles
1
Expert's answer
2020-10-01T17:23:29-0400

Converting radius from miles to km;

1 mile =1.609km

3000=?


="3000\u00d71.609km" =4827km

"r=4827km"


Arch length ="\frac{\theta} {360}2\pi r"


="\frac{7}{360} ×2× \frac{22}{7}×4827"


=589.97km


b) 1nm=1.852km

?nm =589.97km


="\frac{1 ×589.97}{1.852}"


=318.5nm